# SR-Reward: Taking The Path More Traveled

### v1

@@ 5 Experiments
| Ant | 85.71 32.48 | 83.8432.63 | 81.6432.12 | 87.7332.51 | 82.3132.86 |
| Hopper | 108.64 13.88 | 111.289.17 | 108.6914.73 | 110.6410.86 | 109.6911.19 |
| Halfcheetah | 105.55 12.09 | 104.927.61 | 103.7613.64 | 106.235.52 | 106.03 5.20 |
| Walker2d | 73.22 42.41 | 84.5738.40 | 79.6438.73 | 83.0038.69 | 85.57 32.09 |
| Door | 76.76 35.13 | 96.40 21.59 | 98.24 17.44 | 78.14 39.26 | 104.02 8.60 |
| Hammer | 114.21 29.66 | 90.67 52.51 | 111.64 35.91 | 69.90 56.70 | 117.50 21.26 |
| Pen | 101.74 62.93 | 106.67 61.82 | 94.92 64.53 | 107.11 61.78 | 103.52 62.27 |
| Relocate | 92.92 26.67 | 92.57 25.48 | 88.21 32.44 | 80.86 35.57 | 91.58 25.91 |
| PickCube | 93.31 34.67 | — | — | — | 96.67 29.95 |
| StackCube | 53.01 49.71 | — | — | — | 71.03 46.59 |
| TurnFaucet | 19.71 37.87 | — | — | — | 47.23 46.01 |
5.1 Experimental Setup
@@
To evaluate the proposed SR-Reward framework, we integrate it with two distinct offline RL algorithms: f-DVL (Sikchi et al., 2023) and SparseQL (Xu et al., 2023). Both algorithms, which build on foundational concepts from IQL (Kostrikov et al., 2022) and XQL (Garg et al., 2023), have demonstrated enhanced stability and strong performance in offline reinforcement learning settings. In our experiments, we replace the rewards in the offline dataset with those generated by SR-Reward, allowing the reward function to be learned in conjunction with the RL algorithms.
@@
For empirical validation, we utilize the widely-used MuJoCo-based (Todorov et al., 2012) environments for locomotion tasks, and the Adroit hand (Rajeswaran et al., 2018) environments for assessing performance on more realistic, hand-engineered reward tasks. Figure 2 illustrates these environments. We also include three challenging environments from Maniskill2 (Gu et al., 2023), shown in Figure 5. These environments pose greater difficulty than MuJoCo and Adroit due to variable start and goal positions, extended time horizons, and the complexity of manipulating objects with only two fingers. As in real-world applications, the datasets do not include rewards, making them incompatible with traditional offline RL algorithms.
Each agent is trained for one million gradient steps, with three different random seeds used per task. To select the best checkpoint during training, we evaluated the model every 10,000 training steps across 10 episodes. The checkpoint with the highest mean performance from these evaluations was selected. Each selected checkpoint is then run for 100 episodes (300 runs across all seeds). We store the returns from each episode to compute the aggregate results in Table 1.
@@
Hyperparameters remain largely consistent across environments, with key parameters listed in Table 2 (Appendix B). Offline datasets from D4RL (Fu et al., 2020) are used, and we follow their normalization procedures, employing the provided scores for random and expert demonstrators (Appendix D).
@@
1.
How does using SR-Reward + RL compare to BC when provided with sufficient data?
2.
How does using SR-Reward + RL compare to offline RL with access to ground truth rewards?
3.
How does the performance vary with and without the negative sampling strategy?
4.
How is performance affected by reductions in dataset size?
5.
How is performance affected by reductions in dataset quality?
5.2 SR-Reward + RL v.s. BC
Behavioral Cloning (BC) is a straightforward and effective imitation learning algorithm, particularly when large volumes of expert data are available. The D4RL datasets provide over 1,000 expert demonstrations for each MuJoCo task and 5,000 for the Adroit hand environments, which is sufficient for training competitive BC agents. As shown in Table 1, the combination of SR-Reward with RL rivals BC in the MuJoCo, Adroit Hand, and Maniskill2 environments.
Unlike reward-free imitation learning algorithms, offline RL depends on a well-defined reward function. Our results indicate that SR-Reward delivers a meaningful reward signal, enabling offline RL agents to achieve competitive performance. This trend is further illustrated in Figure 6 (Left), where we compare BC and SR-Reward in more challenging environments such as PickCube, StackCube, and TurnFaucet. While both approaches perform similarly on the PickCube task, SR-Reward slightly outperforms BC on the more demanding StackCube and TurnFaucet tasks, which require greater precision for successful completion.
Moreover, our findings highlight SR-Reward as a modular component that can be integrated with any offline RL algorithm. This flexibility opens up new possibilities for applying offline RL to tasks such as robotics, where demonstrations are easier to obtain than designing dense reward functions, or where sparse rewards may not suffice.
5.3 SR-Reward v.s. True Reward
Since SR-Reward serves as a proxy for the true reward, we compare the performance of agents trained with SR-Reward to those trained with the environment’s native reward. As shown in Table 1, in the MuJoCo environments, the performance of SR-Reward combined with RL closely matches that of offline RL agents trained with the true reward, indicating that the dense reward generated by SR-Reward is as informative as the environment-provided reward.
The Adroit hand environments present additional challenges for offline RL due to their narrower distribution of trajectories and the discontinuous, hand-engineered rewards assigned to each task. As with the MuJoCo experiments, we trained the offline RL agents using the hand-engineered rewards from the D4RL datasets, which include a combination of sub-rewards and thresholds. These discontinuous rewards highlight the complexities of manually designing reward functions and further emphasize the benefits of using SR-Reward.
As illustrated in Table 1, agents trained with SR-Reward perform on par with those trained with the environment’s dense reward. This underscores SR-Reward’s effectiveness, particularly in cases where only sparse rewards are available or where hand-engineered reward functions fail to provide a sufficiently informative reward signal.
5.4 Effect of Negative Sampling
The negative sampling strategy is designed to reduce reward values for out-of-distribution state-action pairs. To evaluate its effectiveness, we train SR-Reward both with and without the negative sampling strategy and compare the results in the PickCube, StackCube, and TurnFaucet environments (Figure 5). These environments are notably more challenging than the MuJoCo and Adroit settings due to their longer task horizons, randomized configurations for each run, and the complex dynamics of manipulating objects with only two fingers. The difficulty escalates progressively from relocating an object in PickCube, to stacking objects in StackCube, and finally to turning the faucet in TurnFaucet, which requires continuous contact throughout the movement.
@@ 5.4 Effect of Negative Sampling
As shown in Figure 6 (Right), the use of the negative sampling strategy leads to improved performance across all tasks. Additionally, we compare this approach against the linear decay strategy proposed by Luo et al. (2020), applied to SR-Reward (rather than the value function). Our results demonstrate that the exponential decay kernel used in our negative sampling strategy yields better performance. While the negative sampling strategy consistently enhances results across tasks, the magnitude of its contribution varies depending on the task.
5.5 Data Size Ablation
To investigate the impact of data size on SR-Reward, we trained each algorithm using varying numbers of expert trajectories from the D4RL dataset. We used sparseQL as our offline RL algorithm. Specifically, we evaluated performance using [10, 50, 100, 500, 1000] demonstrations for MuJoCo and [10, 50, 100, 500, 1000, 5000] demonstrations for Adroit hand environments. As shown in Figure 7, agents trained with true reward do not significantly outperform those trained with SR-Reward across different data sizes in all MuJoCo environments.
It’s important to note that MuJoCo environments provide dense and continuous rewards based on the agents’ velocity along the X-axis. The straightforward nature of these rewards presents a challenge for any reward-learning algorithm attempting to outperform them. Agents trained with SR-Reward achieve similar returns in most Adroit hand environments than those trained with the true reward, indicating that SR-Reward can also learn an informative reward function comparable to the hand-engineered rewards offered by the environment.
It is worth mentioning that the high variance observed in the results is indicative of the mixed data quality present in the datasets. Figure 9 in Appendix C illustrates the return distribution of expert datasets across all MuJoCo and Adroit environments. Notably, environments such as Walker2d, Ant and Pen contain a significant number of suboptimal demonstrations, even within their expert datasets. These suboptimal demonstrations contribute to the high variance observed in our results. This observation is consistent with results of Fujimoto et al. (2024); Fujimoto & Gu (2021); Hepburn & Montana (2023), where broadening the distribution of returns in a dataset by mixing datasets of different quality (medium-replay, medium-expert) increases the variance in the results for their offline baseline algorithms.
5.6 Data Quality Ablation
@@ 7 Limitations and Future Work
Our experiments focused exclusively on state-based demonstrations. Extending these methods to visual domains is possible by substituting the encoder in Figure 1 with one capable of extracting meaningful representations from image data. Designing an effective visual encoder introduces a new set of engineering challenges that must be addressed when expanding to visual domains. In addition, a theoretical investigation regarding the convergence properties of using SR as a reward can provide more support for our claim, however, in this paper, we have focused on supporting our claims using empirical results.
@@
One limitation of SR-Reward is the assumed availability of a dataset of optimal trajectories. Although our empirical results indicate a degree of robustness when combining optimal and sub-optimal datasets (Figure 8), the presence of sub-optimal demonstrations can negatively impact SR-Reward since the training process treats optimal and sub-optimal demonstrations equally. Enhancing the ability to control the influence of demonstrations based on their quality could lead to higher-quality rewards and more data-efficient learning, offering a promising direction for future research.
@@
Broader Impact Statement
The ability to learn from demonstrations enables users without technical knowledge to program agents such as industrial or household robots. This technology can have great potential for automating tasks where the industry is facing a labor shortage or where the safety of humans is of concern. On the other hand, such technology can accelerate the loss of jobs due to automation, a trend that raises concerns for many. So far automation in the physical world especially outside of repetitive motions of industrial robots has been limited, however, this can change through further development and deployment of systems that can easily and flexibly learn from a handful of demonstrations.
Appendix A Occupancy Measure and Successor Representations
@@ Appendix B Hyperparameters
Appendix C Dataset Return Distributions
Here we plot the histogram of the return distribution of MuJoCo and Adroit Hand expert datasets provided by D4RL. We can see a correlation between the variance in our results (Figure 7) and the broadness of the return distributions for each environment. The plots show that the "expert" datasets offered by D4RL include sub-optimal trajectories.
To further explore the relationship between dataset diversity and result variability, we plot the coefficient of variation (standard deviation divided by the mean) for both the dataset returns and the returns from the trained SR-Reward models. Figure 10 displays this plot, with environments sorted by decreasing coefficient of variation. The plot reveals that SR-Reward models trained on datasets with lower return variability tend to exhibit lower variation in their returns as well.
To further investigate this relationship, we utilized our Toy Maze environment to generate a small synthetic dataset comprising high-performing trajectories with minimal variation. This dataset consists of 8 expert demonstrations. Figure 11 presents a histogram of the returns, highlighting the low variation in returns. In this environment, sparse rewards are given: -0.001 for each step and 1.0 for reaching the goal. Table 3 summarizes the statistics, comparing the return variations within the dataset to the variations in returns from the SR-Reward + RL model trained on it. These results align with our earlier observations, where lower variation in dataset returns corresponds to lower performance variability in models using SR-Reward as the reward function.
| Source | Mean | Std. Dev. | Coefficient of Variations |
| Dataset | 0.8750 | 0.0020 | 0.0023 |
| SR-Reward + RL (SparseQL) | 0.8799 | 0.0037 | 0.0042 |
| True Reward + RL (SparseQL) | 0.8919 | 0.0008 | 0.0009 |
| BC | 0.8874 | 0.0015 | 0.0016 |
Appendix D D4RL Return Normalization
@@
For Maniskill2 environments (PickCube and StackCube) minimum score is considered 0 when the task is not solved and the maximum score is:

### v2

@@ 4 Experimental Evaluation
| Ant | 86.33 3.91 | 84.354.34 | 82.754.92 | 86.645.14 | 82.257.09 |
| Hopper | 108.73 4.39 | 110.741.14 | 108.253.16 | 110.322.69 | 109.590.52 |
| Halfcheetah | 104.87 1.55 | 104.700.36 | 103.801.99 | 106.231.38 | 105.96 0.82 |
| Walker2d | 71.39 14.95 | 84.4114.49 | 78.485.52 | 84.0213.75 | 74.48 7.85 |
| Door | 76.93 21.81 | 96.97 4.70 | 98.40 3.33 | 78.28 23.98 | 104.10 1.57 |
| Hammer | 114.30 4.59 | 90.53 10.28 | 111.52 9.74 | 71.32 18.50 | 117.17 3.71 |
| Pen | 104.95 5.70 | 109.67 10.06 | 97.79 5.15 | 108.75 3.91 | 105.00 3.57 |
| Relocate | 93.20 4.15 | 92.36 4.35 | 88.22 7.30 | 80.69 8.15 | 90.91 6.99 |
| PickCube | 91.79 4.23 | — | 88.81 5.72 | — | 96.70 2.21 |
| StackCube | 51.41 3.97 | — | 70.71 12.55 | — | 71.03 18.41 |
| TurnFaucet | 20.26 1.27 | — | 35.44 6.83 | — | 51.64 9.41 |
4.1 Setup
@@
To evaluate the proposed reward module, we integrate it with two distinct offline RL algorithms: f-DVL (Sikchi et al., 2023) and SparseQL (Xu et al., 2023). Both algorithms, which build on foundational concepts from IQL (Kostrikov et al., 2022) and XQL (Garg et al., 2023), have demonstrated enhanced stability and strong performance in offline reinforcement learning settings. In our experiments, we replace the rewards in the offline dataset with those generated by SR-Reward, allowing the reward function to be learned in conjunction with the RL algorithms.
@@
For empirical validation, we utilize the widely-used MuJoCo-based (Todorov et al., 2012) environments for locomotion tasks, and the Adroit hand (Rajeswaran et al., 2018) environments for assessing performance on more realistic, hand-engineered rewards. Figure 3 illustrates these environments. We also include three challenging environments from Maniskill2 (Gu et al., 2023), shown in Figure 6. These environments pose greater difficulty than MuJoCo and Adroit due to variable start and goal positions, extended time horizons, and the complexity of manipulating objects with only two fingers. As in real-world applications, these datasets do not include rewards, prohibiting the use of offline RL algorithms. We use SR-Reward in combination with SparseQL as the RL agent, for Maniskill2 environments as well as in ablation studies for data size, data quality, and negative sampling (Appendix A, B, and G).
We use state information from the environment and datasets, and hence do not use an image encoder for our experiments. Each agent is trained for one million gradient steps (two million gradient steps for ManiSkill environments), with five different random seeds used per task. For each training run (across five random seeds), we save the checkpoint that achieves the highest mean return over 25 evaluation rollouts, measured every 10,000 steps during training. After selecting the best checkpoint for each seed, we then evaluate it on 50 fresh rollouts and report the mean and standard deviation across the five seeds (one mean return per seed). Table 1 shows the mean return and standard deviation over five seeds for different environments.
@@
Hyperparameters remain largely consistent across environments, with key parameters listed in Table 2 (Appendix D). Further discussion on the hyperparameters of the negative sampling strategy and how to choose them is presented in Appendix G. We use the offline datasets from D4RL (Fu et al., 2020), and follow their normalization procedures, employing the provided scores for random and expert demonstrators (Appendix E).
@@
4.2 Question 1: Can SR-Reward replace the true reward signal?
To address this question, we use the true reward signal from the environment as a baseline and compare the performance of offline RL algorithms (f-DVL and sparseQL) using the true reward against the reward signal generated by SR-Reward.
The MuJoCo and Adroit Hand environments differ significantly in the complexity of their reward functions. In MuJoCo environments, the objective is relatively simple: the agent must move to the right as quickly as possible. A straightforward reward function based on the agent’s velocity suffices for such tasks. In contrast, Adroit Hand environments present more intricate challenges, such as rotating a pen with one hand or hammering a nail. These tasks require dense reward functions, which are constructed using multiple carefully designed sub-rewards and thresholding.
Table 1 summarizes the performance of the offline algorithms using SR-Reward and the true reward. For the MuJoCo environments, the performance achieved with SR-Reward closely matches that of the true reward, demonstrating that the dense reward generated by SR-Reward is as informative as the environment-provided reward. Notably, in the Adroit Hand environments, SR-Reward yields similar or even higher performance than the true reward. This result highlights the inherent challenges of manually engineering reward functions for complex tasks and underscores the advantages of using SR-Reward, particularly for scenarios where crafting a reward function is not straightforward.
Furthermore, Figure 8 and Figure 9 present ablation studies on dataset size and quality, demonstrating that SR-Reward can produce an informative reward signal and sustain competitive performance even under limited or degraded data conditions. While overall performance decreases as the dataset size is reduced or the data quality deteriorates, this trend affects all algorithms similarly. Notably, RL agents (SparseQL) trained with SR-Reward maintain performance levels comparable to those trained with the true environment reward or to behavioral cloning agents, even when learning from suboptimal data. (See Appendix A and Appendix B for further discussion.)
4.3 Question2: How does SR-Reward + offline RL perform compared to BC?
Behavioral Cloning (BC) is a straightforward and effective method in scenarios where no reward function exists, but ample demonstrations are available. To highlight the advantages of using offline RL equipped with SR-Reward, we compare its performance against BC across MuJoCo, Adroit Hand, and more complex ManiSkill2 robot manipulation tasks. In the latter, no true reward signal is available in the datasets, and the offline RL algorithms rely solely on SR-Reward for reward generation.
As shown in Table 1, combining SR-Reward with offline RL performs on par with BC for MuJoCo and Adroit Hand environments. However, as task complexity increases, the performance gap widens. This trend is particularly evident in the ManiSkill2 environments (Figure 7 (Left)), where BC closely matches the performance of SR-Reward-equipped sparseQL for the relatively simple PickCube task but falls behind in the more demanding StackCube and TurnFaucet scenarios, which require higher precision for successful completion. Similar results hold when combining f-DVL with SR-Reward, although BC shows slightly better performance on a the simpler PickCube task but falls behind on StackCube and TurnFaucet.
We have limited our experiments to two offline RL algorithms (f-DVL and SparseQL), however, our findings underscore the versatility of SR-Reward as a modular component that can, in principle, be compatible with many offline RL algorithms by replacing the reward function. This flexibility makes it well-suited for applications in robotics and other domains where obtaining demonstrations is more feasible than designing dense reward functions, and where RL algorithms have the potential to surpass the limitations of BC.
4.4 Question3: Is negative sampling strategy a necessary component of SR-Reward?
A narrow subset of the overall state-action space covered by the demonstrations can lead function approximation networks to predict erroneously high values when queried for out-of-distribution data. Intuitively, the negative sampling strategy presented in this paper can be beneficial in increasing the robustness of the networks used in SR-Reward by lowering the resulting reward values for state-action pairs outside of demonstrations. To evaluate its effectiveness, we train SR-Reward both with and without the negative sampling strategy and compare the results in the PickCube, StackCube, and TurnFaucet environments (Figure 6).
@@ 4.4 Question3: Is negative sampling strategy a necessary component of SR-Reward?
As shown in Figure 7 (Right), the use of the negative sampling strategy leads to similar or improved performance across all tasks. Additionally, we benchmark this approach against the linear decay strategy proposed by Luo et al. (2020), applied to SR-Reward (rather than the value function). Our results demonstrate that the exponential decay kernel used in our negative sampling strategy yields better performance. While the negative sampling strategy consistently enhances results across tasks, the magnitude of its contribution varies depending on the task.
5 Related Work
The growing need to fine-tune large language models has led to a surge of research focused on learning reward functions from human feedback and using them for model optimization (Christiano et al., 2023; Rafailov et al., 2023; Ethayarajh et al., 2024). These approaches typically assume a Bradley-Terry model (Bradley & Terry, 1952) to represent the reward function and rely on datasets containing both positive and negative feedback. In contrast, our work focuses on learning a reward function solely from expert demonstrations, without requiring any form of negative feedback. There is also a series of works that modify the existing environment reward to gain improvements during training. Vieillard et al. (2020) suggests adding of the policy that is being learned to the reward in temporal difference (TD) learning. The authors argue that the logarithm of the policy is a strong learning signal as it is available even in a sparse reward setting and since its value is close to zero for optimal actions under optimal policy this does not conflict with the optimal control objective.
@@ 7 Limitations and Future Work
SR-Reward assumes the availability of a dataset of optimal trajectories. Although our empirical results indicate a degree of robustness when combining optimal and sub-optimal datasets (Figure 9), the presence of sub-optimal demonstrations can negatively impact SR-Reward since the training process treats optimal and sub-optimal demonstrations equally. Enhancing the ability to control the influence of demonstrations based on their quality could lead to higher-quality rewards and more data-efficient learning, offering a promising direction for future research.
@@
Finally, a theoretical investigation regarding the convergence properties of using SR as a reward or the efficacy of using L2 Norm for converting the SR vector to a scalar reward can provide more support for our claim, however, in this paper, we have focused on supporting our claims using empirical results.
Appendix A Data-Size Ablation
To investigate the impact of data size on SR-Reward, we trained each algorithm using varying numbers of expert trajectories from the D4RL dataset. We used sparseQL as our offline RL algorithm. Specifically, we evaluated performance using [10, 50, 100, 500] demonstrations for MuJoCo and Adroit hand environments. As shown in Figure 8, agents trained with true reward do not significantly outperform those trained with SR-Reward across different data sizes in all MuJoCo environments.
Appendix B Data-Quality Ablation
Appendix C Occupancy Measure and Successor Representations
@@ Appendix D Hyperparameters
Appendix E D4RL Return Normalization
@@
For Maniskill2 environments (PickCube, StackCube, TurnFaucet), the minimum score is considered 0 when the task is not solved, and the maximum score is:
@@
Appendix F Policy Comparison in the Action Space
We visualize the actions proposed by two policies on the Adroit Door task to assess their behavioral similarity. Specifically, we compare two models trained using the f-DVL algorithm: one trained with the true reward provided by the environment, and the other trained with our learned SR-Reward. Since both models achieve comparable mean returns (see Table 1), we examine whether they also produce similar actions, which would indicate that SR-Reward leads to a policy behaviorally close to that of the true reward.
Figure 10 illustrates the actions suggested by each policy when evaluated on the same observations drawn from 50 trajectories in the offline dataset. Each subplot corresponds to one dimension of the action space. The alignment between the two sets of actions across all dimensions suggests that the policies behave similarly—supporting the idea that policies trained using SR-Reward behave similarly to the ones trained using the true reward.
Appendix G Negative Sampling Hyperparameters
The negative sampling method introduced in Section 3.3 is governed by two hyperparameters: and . The parameter controls the amount of noise added to expert state-action pairs to generate negative samples, while determines the width of the Gaussian penalty kernel used to measure the similarity between expert and negative samples in the latent space. To study the sensitivity of our method to these hyperparameters, we trained a series of SparseQL + SR-Reward models on the StackCube environment using a grid of different and values. Figure 11 presents the normalized returns across these settings. The red circle marks the hyperparameter configuration used in our main experiments (Section 4).
Overall, high values of tend to degrade performance, likely because the resulting negative samples become out-of-distribution with respect to the demonstration data, which can negatively affect training. A similar trend is observed for large values of : when the penalty kernel is too wide, both expert and perturbed samples receive similar rewards, making it harder to distinguish optimal from suboptimal behaviors, ultimately harming performance. As a practical guideline, we set to approximately the median of the standard deviations across all observation and action dimensions in the demonstration data. Figure 12 shows the distribution of values across these dimensions, with the red line marking the median standard deviation. This choice typically yields a noise level that reflects the natural variability present in the expert data. For , we choose a value between 3 to 7 times the selected . This ensures that the reward function remains sensitive enough to distinguish perturbed samples from expert ones—penalizing them appropriately—while not being so narrow as to overly suppress reward values for slightly perturbed but still useful samples.
Appendix H Online Reinforcement Learning
SR-Reward can be trained independently of the reinforcement learning (RL) algorithm using only the offline dataset. In this section, we address the question of whether SR-Reward, once trained offline, can be effectively used as a reward function for online reinforcement learning. Figure 13 compares the performance of TD3 trained using the environment’s true reward versus using SR-Reward trained on an offline dataset of demonstrations. The TD3 agent was trained on the HalfCheetah environment for 1 million gradient steps using online interactions.
The results highlight a key limitation of SR-Reward in online settings. While incorporating negative sampling improves its robustness to some extent, there is still a notable drop in performance when SR-Reward is used in place of the environment’s true reward. This outcome is expected: since SR-Reward is trained solely on expert demonstrations, it does not generalize well to parts of the state-action space that lie far from the distribution of the offline dataset. As the online agent explores new regions during training, SR-Reward may assign incorrect values to unfamiliar state-action pairs, potentially misleading the agent and resulting in suboptimal behavior.